### v2

# Unipotent normal subgroups of algebraic groups

Abstract.
Let $G$ be an affine algebraic group scheme over a field $k$. We show there exists a unipotent normal subgroup of $G$ which contains all other such subgroups; we call it the restricted unipotent radical $\operatorname{Rad}_{u}(G)$ of $G$. We investigate some properties of $\operatorname{Rad}_{u}(G)$, and study those $G$ for which $\operatorname{Rad}_{u}(G)$ is trivial. In particular, we relate these notions to their well-known analogues for smooth connected affine $k$-groups.
footnotetext: 2020 Mathematics Subject Classification. Primary 20G15; Secondary 20G07.
1. Introduction
The notion of unipotency plays a key role in the theory of algebraic groups over a field. An algebraic group over a field is unipotent if it embeds in the algebraic group of upper unitriangular matrices of $\operatorname{GL}_{n}$ for some integer $n\geq 1$.
Over an algebraically closed field, a reductive group is a smooth connected affine group which contains no non-trivial smooth connected unipotent normal subgroups. The theory of reductive groups dates back (at least) to the 1950s; with foundational works by Chevalley and Borel. Reductive groups have a wide range of applications, for example to the Langlands program.
Pseudo-reductive groups are an important generalisation of reductive groups, which are relevant only when our underlying field is imperfect. These were first introduced by Borel-Tits in 1978 [BoTi], and have recently received a lot of attention due to the influential work by Conrad-Gabber-Prasad [CGP] in the 2010s. Notably, in 2012 Conrad [Co] used pseudo-reductive groups to prove some finiteness theorems for algebraic groups over global function fields.
The notions of reductive and pseudo-reductive groups make sense due to the following simple fact. Any affine algebraic group $G$ over a field contains a largest smooth connected unipotent normal subgroup; this is called the unipotent radical $\mathscr{R}_{u}(G)$ of $G$.
We will show that an analogue of this fact holds without any smoothness or connectedness assumptions. More precisely, we prove the following:
Theorem 1.1.
Let $k$ be a field. Let $G$ be an affine algebraic $k$-group. There exists a largest unipotent normal subgroup $\operatorname{Rad}_{u}(G)$ of $G$ (that is, a unipotent normal subgroup of $G$ which contains all other unipotent normal subgroups of $G$). The formation of $\operatorname{Rad}_{u}(G)$ commutes with base change by separable algebraic field extensions.
Henceforth we use the following setup. Let $k$ be an arbitrary field, say of characteristic $p\geq 0$. Let $\overline{k}$ be an algebraic closure of $k$, and let $k^{\operatorname{s}}$ be the separable closure of $k$ in $\overline{k}$. Let $G$ be an affine algebraic $k$-group (i.e. an affine group scheme of finite type over $k$).
Let $\operatorname{Rad}_{u}(G)$ be the largest unipotent normal subgroup of $G$ as in Theorem 1.1; we call it the restricted unipotent radical of $G$. In analogue with the notion of pseudo-reductivity, we say that $G$ is pseudo-$p$-reductive if $\operatorname{Rad}_{u}(G)$ is trivial. Similarly, $G$ is $p$-reductive if $\operatorname{Rad}_{u}(G_{\overline{k}})$ is trivial. We will see that there are parallels between the behaviour of such groups and their smooth connected analogues. However, certain complexities arise when working with $\operatorname{Rad}_{u}(G)$ that one does not see with $\mathscr{R}_{u}(G)$.
An example of a commutative infinitesimal $k$-group that is pseudo-$p$-reductive but not $p$-reductive is the $k$-group NN constructed in Example 3.3.
Next, we find sufficient conditions on $G$ under which $\operatorname{Rad}_{u}(G)$ contains all unipotent subgroups of $G$ (not just all unipotent normal subgroups). Of course this holds if $G$ is commutative.
Theorem 1.2.
Let $G$ be an affine algebraic $k$-group. Suppose at least one of the following holds:
(a) $G$ is trigonalisable over $\overline{k}$;
(b) $G\cong\overline{G}/Z(\overline{G})$, for some nilpotent connected algebraic $k$-group $\overline{G}$.
Then $\operatorname{Rad}_{u}(G)$ contains all unipotent subgroups of $G$.
For example, an affine algebraic $k$-group $G$ is trigonalisable over $\overline{k}$ if it is commutative, or if it is smooth, connected and solvable by the Lie-Kolchin theorem [Mi, Thm. 16.30].
An example of a nilpotent infinitesimal $k$-group which does not contain a largest unipotent subgroup is the Frobenius kernel $G=(\operatorname{SL}_{2})_{1}$ of $\operatorname{SL}_{2}$ in characteristic 2. To see this, let $T$ be a maximal torus of $\operatorname{SL}_{2}$ with corresponding root groups $U^{+}$ and $U^{-}$. The center $Z(G)$ of $G$ equals $T_{1}$, which is of multiplicative type, so $G$ is nilpotent but not unipotent. However $G$ is generated by $(U^{+})_{1}$ and $(U^{-})_{1}$, both of which are unipotent.
If $G$ is an affine algebraic $k$-group such that all unipotent subgroups of $G$ are contained in $\operatorname{Rad}_{u}(G)$, then $G/\operatorname{Rad}_{u}(G)$ contains no non-trivial unipotent subgroups. This motivates the following question: which affine algebraic $k$-groups contain no non-trivial unipotent subgroups?
Clearly any algebraic $k$-group of multiplicative type satisfies this property; for example tori, or the $n$’th roots of unity $\mu_{n}$ for any integer $n$. A wider class of examples may be constructed as follows: let $k^{\prime}/k$ be a finite field extension, and let $Z^{\prime}$ be an algebraic $k^{\prime}$-group of multiplicative type. Then the Weil restriction $R_{k^{\prime}/k}(Z^{\prime})$ is commutative and contains no non-trivial unipotent subgroups.
It turns out that in many important situations, affine algebraic $k$-groups which contain no non-trivial unipotent subgroups are automatically commutative (in the language of [CP, §1.4], such groups are called tame). More precisely, the following is true.
Theorem 1.3.
Let $G$ be a connected affine algebraic $k$-group. Suppose $G$ contains no non-trivial unipotent subgroups. Then $G$ is commutative in each of the following cases:
(a) $k$ is algebraically closed;
(b) $k$ is perfect and $G$ is solvable;
(c) $G$ is trigonalisable over $\overline{k}$.
We illustrate Theorems 1.2 and 1.3 with several examples of affine algebraic $k$-groups $G$ which contain no non-trivial unipotent subgroups, and yet are neither nilpotent nor trigonalisable over $\overline{k}$. Firstly, over any field $k$, take any perfect finite $k$-group $G$ with order coprime to $p$. Secondly, if $k=\mathbb{R}$, take any anisotropic reductive $k$-group $G$ that is not a torus (e.g. $\operatorname{SU}_{2}$). Finally, over an imperfect field $k$, we construct a solvable infinitesimal such $G$ in Example 3.3.
We now present some consequences of Theorems 1.2 and 1.3.
We need the following definition; which is taken from [CGP, (1.3.1)] (it is most useful for smooth $G$, but makes sense for arbitrary $G$). Let $k^{\prime}/k$ be the minimal field of definition for $\mathscr{R}_{u}(\smash{G_{\overline{k}}})$, and let $\pi^{\prime}:G_{k^{\prime}}\rightarrow G_{k^{\prime}}/\mathscr{R}_{u}(G_{k^{\prime}})=:G^{\prime}$ be the natural projection. Let
$i_{G}:G\rightarrow R_{k^{\prime}/k}(G^{\prime})$ (1)
be the $k$-homomorphism associated to $\pi^{\prime}$ under the adjunction of extension of scalars and Weil restriction (refer to [CGP, (A.5.1)] for this adjunction).
Corollary 1.4.
[…]
Corollary 1.5.
Let $G$ be a connected algebraic $k$-group which satisfies either condition (a) or (b) of Theorem 1.2. Then $G/\operatorname{Rad}_{u}(G)$ embeds in a Weil restriction $R_{k^{\prime}/k}(T^{\prime})$, for some purely inseparable finite field extension $k^{\prime}/k$ and $k^{\prime}$-torus $T^{\prime}$.
Corollary 1.6.
Assume $p>0$. Let $G$ be a connected affine algebraic $k$-group. Suppose the Frobenius kernel $G_{1}$ of $G$ is of multiplicative type. Then $G$ is of multiplicative type.
Corollary 1.7.
Assume $p>0$. Let $G$ be a connected affine algebraic $k$-group. Suppose the Frobenius kernel $G_{1}$ of $G$ is trigonalisable over $\overline{k}$ and $p$-reductive. Then $G$ is of multiplicative type.
Next, we look at the relationship between pseudo-$p$-reductivity and the usual notion of pseudo-reductivity for smooth connected affine $k$-groups.
It is known that, as long as $p>2$, a smooth connected affine $k$-group $G$ is reductive if and only if its Frobenius kernel $G_{1}$ is $p$-reductive; see for instance [Hu, Prop. 11.8]. The following theorem generalises this result.
Theorem 1.8.
[…]
(i) Assume $p>2$. Then $G_{1}$ is pseudo-$p$-reductive if and only if $\ker i_{G}$ is étale. If in addition $G$ is perfect, this condition holds if and only if $G$ is pseudo-reductive of minimal type.
(ii) Assume $p=2$. Then $G_{1}$ is pseudo-$p$-reductive if and only if $\ker i_{G}$ is étale and $(G^{\prime})_{k^{\prime\operatorname{s}}}$ has no direct factor isomorphic to $\mathrm{SO}_{2n+1}$ for some $n\geq 1$.
Corollary 1.9.
Assume $p>0$. Let $G$ be a smooth connected affine $k$-group. Suppose the Frobenius kernel $G_{1}$ of $G$ is pseudo-$p$-reductive. Then $G$ is pseudo-reductive and its $k^{\operatorname{s}}$-root system is reduced.
Despite these similarities, the behaviour of $p$-reductive and pseudo-$p$-reductive $k$-groups differs in several important ways from their smooth connected analogues. Notably, $p$-reductivity is not preserved by taking normal subgroups nor by quotients; again refer to Example 3.3.
2. The restricted unipotent radical
In this section we prove Theorem 1.1. The author thanks Sean Cotner and Timm Peerenboom for some helpful discussions, and the proof of [CP, Prop. 5.1.2] for some ideas.
Recall that $k$ is an arbitrary field. By a downward (resp. upward) directed set, we mean a partially ordered set in which every finite subset has a lower (resp. upper) bound.
Lemma 2.1.
Let $A$ be a $k$-algebra. Let $\{I_{i}\hskip 1.42262pt|\hskip 1.42262pti\in\Lambda\}$ be a (possibly uncountable) set of ideals of $A$ that is downward directed under inclusion (i.e. $\Lambda$ is a downward directed set and, for $i,j\in\Lambda$, if $i\leq j$ then $I_{i}\subseteq I_{j}$). Let $I:=\bigcap_{i\in\Lambda}I_{i}$. Then
$A\otimes_{k}I+I\otimes_{k}A={\bigcap}_{i\in\Lambda}(A\otimes_{k}I_{i}+I_{i}\otimes_{k}A).$
Proof.
We first check that
$A\otimes_{k}I+I\otimes_{k}A={\bigcap}_{i\in\Lambda}(A\otimes_{k}I_{i}+I\otimes_{k}A).$ (2)
[…]
$A\otimes_{k}I+I\otimes_{k}A={\bigcap}_{i,j\in\Lambda}(A\otimes_{k}I_{i}+I_{j}\otimes_{k}A).$ (3)
We then use the fact that $\Lambda$ is a directed set, giving us
${\bigcap}_{i,j\in\Lambda}(A\otimes_{k}I_{i}+I_{j}\otimes_{k}A)={\bigcap}_{l\in\Lambda}(A\otimes_{k}I_{l}+I_{l}\otimes_{k}A).$ (4)
[…]
Now let $X$ be an affine algebraic scheme over $k$. Let $\{Y_{i}\hskip 1.42262pt|\hskip 1.42262pti\in\Lambda\}$ be a set of closed subschemes of $X$. We define the schematic union $\bigcup_{i\in\Lambda}Y_{i}=:Y$ as follows. Let $\xi:\coprod_{i\in\Lambda}Y_{i}\rightarrow X$ be the morphism of $k$-schemes given by inclusion on each component (indeed it exists, as the category of $k$-schemes admits all – possibly uncountable – coproducts). We define $Y$ to be the schematic image of $\xi$, in the sense of [Stacks, 01R7]. Equivalently, $Y$ is the closed subscheme of $X$ whose associated ideal is the intersection of those of all of the $Y_{i}$’s.
In this paper a subgroup scheme (or more simply subgroup) refers to (the equivalence class associated to) a locally closed immersion of group schemes. Following this convention, by [Stacks, 047T], all subgroup schemes of a group scheme over a field $k$ are automatically closed.
Proposition 2.2.
Let $G$ be an affine algebraic $k$-group. Let $\{H_{i}\hskip 1.42262pt|\hskip 1.42262pti\in\Lambda\}$ be a set of (closed) subgroups of $G$ that is upward directed under inclusion. The schematic union $\bigcup_{i\in\Lambda}H_{i}=:H$ is a (closed) subgroup of $G$. If each $H_{i}$ is a normal subgroup of $G$ then $H$ is also a normal subgroup of $G$.
Proof.
By definition of a $k$-group scheme, $G$ comes equipped with the following morphisms of $k$-schemes: $e:1\rightarrow G$ is the unit map, $m:G\times G\rightarrow G$ is multiplication, and $\iota:G\rightarrow G$ is inversion. A subscheme $H$ of $G$ is a subgroup if and only if it is stabilised by inversion, and the restriction $m|_{H\times H}$ factors through the inclusion $H\hookrightarrow G$. It is clear that inversion stabilises $H$, since it is an isomorphism of schemes which stabilises each $H_{i}$, and $H$ is the smallest closed subscheme of $G$ which contains all $H_{i}$’s. We now check multiplication.
Let $G=\operatorname{Spec}A$ and $H_{i}=\operatorname{Spec}A/I_{i}$, for $A$ a finitely generated $k$-algebra and each $I_{i}$ an ideal of $A$. Denote $I:=\bigcap_{i\in\Lambda}I_{i}$. By definition $H=\operatorname{Spec}A/I$. Let $m=\operatorname{Spec}\overline{m}$, where $\overline{m}:A\rightarrow A\otimes_{k}A$ is a $k$-algebra homomorphism.
Fix some $i\in\Lambda$. Since $H_{i}$ is a subgroup scheme of $G$, the map
$A\xrightarrow{\overline{m}}A\otimes_{k}A\rightarrow A/I_{i}\otimes_{k}A/I_{i}=(A\otimes_{k}A)/(A\otimes_{k}I_{i}+I_{i}\otimes_{k}A)$
factors (on the left) through the natural projection $A\rightarrow A/I_{i}$. For $a_{i}\in I_{i}$, this means that
$\overline{m}(a_{i})\in A\otimes_{k}I_{i}+I_{i}\otimes_{k}A\subseteq A\otimes_{k}A.$
Hence, for $a\in I$, we have that
$\overline{m}(a)\in{\bigcap}_{i\in\Lambda}(A\otimes_{k}I_{i}+I_{i}\otimes_{k}A)=A\otimes_{k}I+I\otimes_{k}A$
where the second equality is Lemma 2.1. We deduce that the map
$A\xrightarrow{\overline{m}}A\otimes_{k}A\rightarrow A/I\otimes_{k}A/I=(A\otimes_{k}A)/(A\otimes_{k}I+I\otimes_{k}A)$
factors (on the left) through the natural projection $A\rightarrow A/I$.
So indeed $H$ is a subgroup scheme of $G$. It remains to prove the normality assertion.
We will need the following two facts. Let $\psi:G_{1}\rightarrow G_{2}$ be a homomorphism of affine algebraic $k$-groups, and let $\psi^{*}:\mathcal{O}(G_{2})\rightarrow\mathcal{O}(G_{1})$ be the comorphism (between the coordinate algebras). Firstly, $\psi$ is faithfully flat if and only if $\psi^{*}$ is injective [Mi, Prop. 5.43]. Secondly, $\mathcal{O}(\ker\psi)=\mathcal{O}(G_{1})/\psi^{*}(I_{G_{2}})$; where $I_{G_{2}}=\ker(\mathcal{O}(G_{2})\rightarrow k)$ is the augmentation ideal [Mi, Ch. 1e].
Now recall our previous setup: $G=\operatorname{Spec}A$, $H_{i}=\operatorname{Spec}A/I_{i}$ for each $i\in\Lambda$, and $H=\operatorname{Spec}A/I$ where $I:=\bigcap_{i\in\Lambda}I_{i}$. Suppose each $H_{i}$ is a normal subgroup of $G$. For each $i\in\Lambda$, let $f_{i}:G\rightarrow G/H_{i}=:Q_{i}$ be the natural projection and say $Q_{i}=\operatorname{Spec}B_{i}$. Let $B:=\bigcap_{i\in\Lambda}B_{i}$ and $Q:=\operatorname{Spec}B$. Choose any $i\in\Lambda$, and define $f:G\rightarrow Q$ to be the spectrum of the composite
$B\lhook\joinrel\longrightarrow B_{i}\xrightarrow{\smash{(f_{i})^{*}}}A.$
Observe that $f$ is a faithfully flat homomorphism of affine $k$-groups, since $f^{*}$ is an embedding of Hopf $k$-algebras.
Using the aforementioned two facts, we compute
$f^{*}(I_{Q})$ $=f^{*}({\bigcap}_{i\in\Lambda}\ker(B_{i}\rightarrow k))$
$={\bigcap}_{i\in\Lambda}(f_{i})^{*}(\ker(B_{i}\rightarrow k))$
$={\bigcap}_{i\in\Lambda}I_{i}=I$
Hence $H=\ker f$. This completes the proof. ∎
We now have enough to prove Theorem 1.1.
Proof of Theorem 1.1.
Let $G$ be an affine algebraic $k$-group.
Proof.
Let $U_{1}$ and $U_{2}$ be unipotent normal subgroups of $G$. Since the property of unipotency is preserved by taking quotients and group extensions, $U_{1}U_{2}$ is also a unipotent normal subgroup of $G$. So the unipotent normal subgroups of $G$ form an upward directed set under inclusion; let us call it $\mathcal{U}$.
Let $\operatorname{Rad}_{u}(G):=\bigcup_{U\in\mathcal{U}}U$; the schematic union. By construction $\operatorname{Rad}_{u}(G)$ contains all unipotent normal subgroups of $G$. Applying Proposition 2.2 tells us that $\operatorname{Rad}_{u}(G)$ is itself a normal subgroup of $G$. We next show that $\operatorname{Rad}_{u}(G)$ is unipotent.
Since $\operatorname{Rad}_{u}(G)_{\overline{k}}\subseteq\operatorname{Rad}_{u}(G_{\overline{k}})$, it suffices to assume that $k=\overline{k}$. Consider the (not necessarily Zariski-closed) subgroup $\bigcup_{U\in\mathcal{U}}U(k)$ of the abstract group $G(k)$. Since all elements of $\bigcup_{U\in\mathcal{U}}U(k)$ are unipotent, this implies that $\bigcup_{U\in\mathcal{U}}U(k)$ is strictly upper trigonalisable (by the same argument as for closed subgroups). Hence $\operatorname{Rad}_{u}(G)(k)$ is strictly upper trigonalisable, as it equals the Zariski-closure of $\bigcup_{U\in\mathcal{U}}U(k)$. Note that $\operatorname{Rad}_{u}(G)_{\operatorname{red}}$ is a smooth subgroup of $G$, as $\smash{k=\overline{k}}$. Then applying [SGA3, XVII, Cor. 3.8] tells us that $\operatorname{Rad}_{u}(G)_{\operatorname{red}}$ is unipotent. This takes care of the case where $p=0$ by Cartier’s theorem [Mi, Thm. 3.23], so we can assume that $p>0$.
Let $i\geq 1$ be an integer. Consider the schematic union $\bigcup_{U\in\mathcal{U}}U_{i}$, where $U_{i}$ denotes the $i$’th Frobenius kernel of $U$. Note that $\bigcup_{U\in\mathcal{U}}U_{i}$ is a finite $k$-group, as it is contained in $G_{i}$. Hence, since $\mathcal{U}$ is an upward directed set, there exists some $V\in\mathcal{U}$ such that $V_{i}=\bigcup_{U\in\mathcal{U}}U_{i}$. Consequently, $\bigcup_{U\in\mathcal{U}}U_{i}$ is a unipotent normal subgroup of $G$.
Now choose a sufficiently large integer $t$ such that the chain of subgroups
$\big(({\bigcup}_{U\in\mathcal{U}}U_{i})\cdot\operatorname{Rad}_{u}(G)_{\operatorname{red}}\big)_{i\geq 1}$
stabilises for all $i\geq t$ (this is indeed possible as $\operatorname{Rad}_{u}(G)_{\operatorname{red}}$ is a finite index subgroup of $\operatorname{Rad}_{u}(G)$). Let $H\in\mathcal{U}$. We apply the following key result taken from [SGA3, VIIA, Prop. 8.3] to $H$: there exists a sufficiently large integer $r$ such that $H/H_{r}$ is smooth. Since $k=\overline{k}$, this implies that $H=H_{r}\cdot H_{\operatorname{red}}$. Then
$H\subseteq({\bigcup}_{U\in\mathcal{U}}U_{t})\cdot\operatorname{Rad}_{u}(G)_{\operatorname{red}}.$
Hence $\operatorname{Rad}_{u}(G)=(\bigcup_{U\in\mathcal{U}}U_{t})\cdot\operatorname{Rad}_{u}(G)_{\operatorname{red}}$, since $H\in\mathcal{U}$ was chosen arbitrarily. Then, using [Mi, Lem. 6.41], we deduce that $\operatorname{Rad}_{u}(G)$ is unipotent.
The invariance of $\operatorname{Rad}_{u}(G)$ under base change by separable algebraic field extensions follows from a simple argument of Galois descent. Explicitly, let $k^{\prime}/k$ be a separable algebraic field extension. By uniqueness $\operatorname{Rad}_{u}(G_{k^{\prime}})$ is stable under the action of the Galois group $\operatorname{Gal}(k^{\prime}/k)$, so it descends to $k$. It follows that $\operatorname{Rad}_{u}(G)_{k^{\prime}}=\operatorname{Rad}_{u}(G_{k^{\prime}})$, as the properties of unipotency and normality are invariant under base change. This completes the proof. ∎
3. The largest unipotent subgroup
Recall that $k$ is a field, of characteristic $p\geq 0$.
In this section we investigate the conditions under which an affine algebraic $k$-group admits a largest unipotent subgroup. Along the way we prove Theorems 1.2 and 1.3, as well as Corollaries 1.4, 1.5, 1.6 and 1.7.
Proof of Theorem 1.2.
Let $G$ be an affine algebraic $k$-group (in fact the word affine is superfluous).
Proof.
Proof of (a). Suppose $G$ is trigonalisable over $\overline{k}$.
The formation of the derived group $\mathscr{D}(G)$ commutes with arbitrary base change by [Mi, Cor. 6.19(a)]. Hence $\mathscr{D}(G)$ is unipotent by [Mi, Thm. 16.6], since $G$ is trigonalisable over $\overline{k}$. Consider the extension
$1\rightarrow\mathscr{D}(G)\rightarrow G\xrightarrow{\kappa}G/\mathscr{D}(G)\rightarrow 1.$
Since $G/\mathscr{D}(G)$ is commutative, its restricted unipotent radical $\operatorname{Rad}_{u}(G/\mathscr{D}(G))$ contains all unipotent subgroups of $G/\mathscr{D}(G)$. In other words, $\operatorname{Rad}_{u}(G/\mathscr{D}(G))$ is the largest unipotent subgroup of $G/\mathscr{D}(G)$. Consider the normal subgroup $U:=\kappa^{-1}(\operatorname{Rad}_{u}(G/\mathscr{D}(G)))$ of $G$. Recall that the property of unipotency is preserved by taking quotients and group extensions. Since $\mathscr{D}(G)$ is unipotent, it follows that $U$ is the largest unipotent subgroup of $G$. Hence $\operatorname{Rad}_{u}(G)=U$, by normality of $U$.
Proof of (b). Let $\overline{G}$ be a connected algebraic $k$-group, and let $G=\overline{G}/Z(\overline{G})$. Then its center $Z(G)$ is unipotent by [SGA3, XVII, Lem. 7.3.2]. [N.B. $G$ is automatically affine by [Mi, Cor. 8.11].]
Suppose in addition that $\overline{G}$, and hence $G$, is nilpotent. Let $n$ be the nilpotency class of $G$. We induct on $n$. If $n=1$ then $G$ is commutative, and the result is given by Theorem 1.1. For $n>1$, consider the central extension
$1\rightarrow Z(G)\rightarrow G\xrightarrow{\zeta}G/Z(G)\rightarrow 1.$ (5)
The nilpotency class of $G/Z(G)$ is n1n-1, so by the inductive hypothesis $\operatorname{Rad}_{u}(G/Z(G))$ is the largest unipotent subgroup of $G/Z(G)$. Consider the normal subgroup $V:=\zeta^{-1}(\operatorname{Rad}_{u}(G/Z(G)))$ of $G$. Using the fact that $Z(G)$ is unipotent, a similar argument as in (a) tells us that $\operatorname{Rad}_{u}(G)=V$. ∎
We now move on to the proof of Theorem 1.3. We will repeatedly use the following easy lemma.
Lemma 3.1.
Assume $p>0$. Let $G$ be an infinitesimal $k$-group. Then either $G$ contains a copy of $\alpha_{p}$, or $G$ contains no non-trivial unipotent subgroups.
Proof.
Assume that $G$ contains a non-trivial unipotent subgroup $U$. Since $G$ is infinitesimal, then so is $U$. We can then find a copy of $\alpha_{p}$ in $U$ as the penultimate term of a composition series for $U$; this follows from [SGA3, XVII, Prop. 4.2.1]. ∎
Proof of Theorem 1.3.
Let $G$ be a connected affine algebraic $k$-group which contains no non-trivial unipotent subgroups.
Proof.
Cases (a) and (b). Suppose $k$ is perfect. In this case, showing that $G$ is commutative is equivalent to showing that $G$ is of multiplicative type; this follows from [DG, IV, §3, Thm. 1.1]. Suppose in addition that either $k$ is algebraically closed, or $G$ is solvable.
First, assume that $G$ is smooth. Since $k$ is perfect and $G$ has no non-trivial unipotent subgroups, it is reductive. In the case where $k$ is algebraically closed: $G$ is split and all of its root groups are trivial, so $G$ is a torus. In the case where $G$ is solvable: $G$ is again a torus, as any solvable reductive $k$-group is a torus. So, henceforth, assume that $G$ is not smooth. In particular, $p>0$ by Cartier’s theorem [Mi, Thm. 3.23].
Next, assume that $G$ is infinitesimal of height $1$. The functor $G\mapsto\operatorname{Lie}(G)$ is an equivalence between the category of height $\leq\!1$ $k$-groups and the category of finite-dimensional restricted Lie algebras; see [DG, II, §7, Prop. 4.1]. Under this equivalence, unipotent height $\leq\!1$ $k$-groups correspond to $p$-nilpotent restricted Lie algebras by [SGA3, XVII, Cor. 3.7]. Since $G$ contains no non-trivial unipotent subgroups, $\operatorname{Lie}(G)$ contains no non-trivial $p$-nilpotent elements. We deduce that $\operatorname{Lie}(G)$ – and hence $G$ – is commutative (the case where $k$ is algebraically closed is [FS, §2, Thm. 3.10], the case where $G$ is solvable is [FS, §2, Prop. 3.9]). Then $G$ of multiplicative type by [DG, IV, §3, Thm. 1.1].
Finally, consider the general (non-smooth) case. For any integer $i$, let $G_{i}$ denote the $i$’th Frobenius kernel of $G$. By [SGA3, VIIA, Prop. 8.3] there exists a sufficiently large integer $r$ such that $G/G_{r}$ is smooth. Take $r$ to be minimal with respect to this condition, and consider the subnormal series
$G\supseteq G_{r}\supset G_{r-1}\supset...\supset G_{1}\supset 1.$ (6)
[…]
Next, say $r>1$ and $G=G_{r}$. Then $G$ is an extension of the height 1 $k$-group $G/G_{r-1}$ by the height r1r-1 $k$-group $G_{r-1}$; let us call the natural projection $\rho_{r}:G\rightarrow G/G_{r-1}$. By the inductive hypothesis, $G_{r-1}$ is of multiplicative type. Assume (for a contradiction) that $G/G_{r-1}$ contains a non-trivial unipotent subgroup. Since $G/G_{r-1}$ is infinitesimal, it contains a subgroup $U$ which is isomorphic to $\alpha_{p}$ by Lemma 3.1. Then, since $k$ is perfect and $G_{r-1}$ is of multiplicative type, [SGA3, XVII, Thm. 6.1.1(B)] tells us that the extension
$1\rightarrow G_{r-1}\rightarrow\rho_{r}^{-1}(U)\rightarrow U\rightarrow 1$
splits, which is a contradiction. [N.B. In fact $\rho_{r}^{-1}(U)\cong G_{r-1}\times U$, but we do not need this.] So indeed $G/G_{r-1}$ contains no non-trivial unipotent subgroups. Hence $G/G_{r-1}$ is of multiplicative type by the inductive hypothesis. Then applying [Mi, Thm. 15.39] tells us that $G$ is also of multiplicative type.
Finally, say $G\supsetneq G_{r}$. The argument here is very similar to that of the previous paragraph. We have that $G$ is an extension of the smooth $k$-group $G/G_{r}$ by the infinitesimal $k$-group $G_{r}$. By the inductive hypothesis, $G_{r}$ is of multiplicative type. Assume (for a contradiction) that $G/G_{r}$ contains a non-trivial unipotent subgroup $U$. Then either $U$ is étale or its Frobenius kernel $U_{1}$ is non-trivial, in which case $U$ contains a copy of $\alpha_{p}$ by Lemma 3.1. In either case [SGA3, XVII, Thm. 6.1.1(A,B)] gives us a contradiction (as étale $k$-groups are smooth). Hence $G$ is of multiplicative type, again by [Mi, Thm. 15.39]. In particular, $G$ is commutative. This completes the proof for Cases (a) and (b).
Case (c). Suppose that $G$ is trigonalisable over $\overline{k}$. Consider the derived group $\mathscr{D}(G)$ as defined in [Mi, Def. 6.16]. Since $G$ is affine, by [Mi, Cor. 6.19(a)] the formation of $\mathscr{D}(G)$ commutes with base change by $\overline{k}/k$. Then base changing by $\overline{k}/k$ and applying [Mi, Thm. 16.6] tells us that $\mathscr{D}(G)$ is unipotent. Hence $\mathscr{D}(G)$ is trivial by assumption. That is, $G$ is commutative. This completes the proof. ∎
Proof of Corollary 1.4.
Let $G$ be a solvable smooth connected affine $k$-group.
Proof.
The Lie-Kolchin theorem [Mi, Thm. 16.30] says that $G$ is trigonalisable over $\overline{k}$. Hence $\operatorname{Rad}_{u}(G)$ contains all unipotent subgroups of $G$, by Theorem 1.2(a). We now base change by $\overline{k}/k$. A similar argument reassures us that $\operatorname{Rad}_{u}(G_{\overline{k}})$ contains all unipotent subgroups of $G_{\overline{k}}$. So $G_{\overline{k}}/\operatorname{Rad}_{u}(G_{\overline{k}})$ contains no non-trivial unipotent subgroups. Then applying Theorem 1.3 says that $G_{\overline{k}}/\operatorname{Rad}_{u}(G_{\overline{k}})$ is commutative, and hence it is of multiplicative type by [DG, IV, §3, Thm. 1.1]. Consequently, the extension
$1\rightarrow\operatorname{Rad}_{u}(G_{\overline{k}})\rightarrow G_{\overline{k}}\rightarrow G_{\overline{k}}/\operatorname{Rad}_{u}(G_{\overline{k}})\rightarrow 1$
splits by [SGA3, XVII, Thm. 5.1.1]. In particular $\operatorname{Rad}_{u}(G_{\overline{k}})$ is smooth and connected, so $\operatorname{Rad}_{u}(G_{\overline{k}})=\mathscr{R}_{u}(G_{\overline{k}})$. By construction $\ker i_{G}$ is the largest subgroup of $G$ satisfying the property that $(\ker i_{G})_{\overline{k}}\subseteq\mathscr{R}_{u}(G_{\overline{k}})$. Hence $\ker i_{G}=\operatorname{Rad}_{u}(G)$. ∎
Proof of Corollary 1.5.
We prove a slightly more general result which implies Corollary 1.5.
Let $(P)$ be a property of algebraic group schemes over a field which is preserved by base change by arbitrary (algebraic) field extensions and such that, for any connected affine algebraic $k$-group $G$ with property $(P)$, all unipotent subgroups of $G$ are contained in $\operatorname{Rad}_{u}(G)$.
Proposition 3.2.
Let $G$ be a connected affine algebraic $k$-group $G$ with property $(P)$. Then $G/\operatorname{Rad}_{u}(G)$ embeds in a Weil restriction $R_{k^{\prime}/k}(T^{\prime})$; for some purely inseparable finite field extension $k^{\prime}/k$ and $k^{\prime}$-torus $T^{\prime}$.
Proof.
By assumption $G_{\overline{k}}$ also satisfies $(P)$. Consider the restricted unipotent radical $\operatorname{Rad}_{u}(G_{\overline{k}})$ of $G_{\overline{k}}$. Let $k^{\prime}/k$ be the minimal field of definition for $\operatorname{Rad}_{u}(G_{\overline{k}})$; it is purely inseparable as the formation of $\operatorname{Rad}_{u}(G)$ commutes with separable field extensions, and it is finite by a similar argument as in [CGP, Prop. 1.1.9(b)].
Consider the natural projection $\rho^{\prime}:G_{k^{\prime}}\rightarrow G_{k^{\prime}}/\operatorname{Rad}_{u}(G_{k^{\prime}})=:G^{\prime}$. Since $\operatorname{Rad}_{u}(G_{k^{\prime}})$ is the largest unipotent subgroup of $G_{k^{\prime}}$ and as the property of unipotency is preserved by group extensions, it follows that $G^{\prime}$ contains no non-trivial unipotent subgroups. Combining Theorem 1.3(a) with [DG, IV, §3, Thm. 1.1] implies that $G^{\prime}$ is of multiplicative type (as it is a $k^{\prime}$-descent of a multiplicative type $\overline{k}$-group). Hence $G^{\prime}$ is contained in some $k^{\prime}$-torus $T^{\prime}$. Since the Weil restriction functor preserves embeddings, we have that $R_{k^{\prime}/k}(G^{\prime})$ embeds in $R_{k^{\prime}/k}(T^{\prime})$.
Let $\iota:G\rightarrow R_{k^{\prime}/k}(G^{\prime})$ be the map associated to $\rho^{\prime}$ under the adjunction of extension of scalars and Weil restriction. The universal property of the counit of this adjunction implies that $(\ker\iota)_{k^{\prime}}\subseteq\ker\rho^{\prime}=\operatorname{Rad}_{u}(G_{k^{\prime}})$; in particular, $\ker\iota$ is unipotent. So $\ker\iota\subseteq\operatorname{Rad}_{u}(G)$. In summary, we have embeddings
$G/\operatorname{Rad}_{u}(G)\hookrightarrow G/\ker\iota\hookrightarrow R_{k^{\prime}/k}(G^{\prime})\hookrightarrow R_{k^{\prime}/k}(T^{\prime}).\qed$
Clearly properties (a) and (b) of Theorem 1.2 are preserved by base change by arbitrary field extensions. Then Corollary 1.5 follows immediately from combining Theorem 1.2 with Proposition 3.2.
Proof of Corollary 1.6.
Assume $p>0$. Let $G$ be a connected affine algebraic $k$-group. Suppose the Frobenius kernel $G_{1}$ of $G$ is of multiplicative type.
Proof.
The formation of the Frobenius map commutes with arbitrary base change, so without loss of generality we can assume $k=\overline{k}$. If $G$ contains a subgroup $U$ that is isomorphic to $\alpha_{p}$ then $U$ is contained in $G_{1}$, which contradicts the fact that $G_{1}$ is of multiplicative type. So $G$ does not contain a copy of $\alpha_{p}$.
First assume that $G$ is smooth. Since $k$ is algebraically closed and $G$ does not contain a copy of $\alpha_{p}$, we see that $G$ is reductive and all of its root groups are trivial. Consequently, $G$ is a torus; which is of multiplicative type.
Next, assume that $G$ is not smooth. By [SGA3, VIIA, Prop. 8.3] there exists a sufficiently large integer $r$ such that $G/G_{r}$ is smooth; where $G_{r}$ denotes the $r$’th Frobenius kernel of $G$. We have a short exact sequence
$1\rightarrow G_{r}\rightarrow G\xrightarrow{f}G/G_{r}\rightarrow 1.$ (7)
Consider the first Frobenius kernel $(G/G_{r})_{1}$ of $G/G_{r}$, and its preimage $f^{-1}((G/G_{r})_{1})=:H$ in $G$.
[…]
Proof of Corollary 1.7.
Assume $p>0$. Let $G$ be a connected affine algebraic $k$-group. Suppose the Frobenius kernel $G_{1}$ of $G$ is $\overline{k}$-trigonalisable and $p$-reductive.
Proof.
Since $G_{1}$ is $\overline{k}$-trigonalisable and $\operatorname{Rad}_{u}(G_{1})$ is trivial, Theorem 1.2(a) says that $G_{1}$ contains no non-trivial unipotent subgroups. Hence $G_{1}$ is commutative by Theorem 1.3(c). Extending scalars by $\overline{k}/k$ and applying [DG, IV, §3, Thm. 1.1] tells us that $G_{1}$ is of multiplicative type, as it is $p$-reductive. Hence $G$ is of multiplicative type by Corollary 1.6. ∎
In the following example we construct a solvable infinitesimal $k$-group $G$ which is not nilpotent, and contains no non-trivial unipotent subgroups. It is a consequence of Theorem 1.3 that $G$ is not trigonalisable over $\overline{k}$.
Example 3.3.
Let $k$ be an imperfect field of characteristic $p>0$, and let $a\in k\setminus k^{p}$. Let $\mathfrak{g}$ be the 3-dimensional restricted Lie algebra with generators $X,Y,Z$ satisfying $[X,Y]=[X,Z]=0$, $[Z,Y]=Y$, $X^{(p)}=X$, $Y^{(p)}=aX$, $Z^{(p)}=Z$. Explicitly, we have a faithful restricted representation $\mathfrak{g}\rightarrow\mathfrak{gl}_{p}$ given by $X\mapsto\operatorname{diag}(1,1,...,1)$, $Y\mapsto(y_{ij})_{i,j}$ where $y_{1p}=a$, $y_{ij}=1$ if j=i1j=i-1 and $y_{ij}=0$ otherwise, and $Z\mapsto\operatorname{diag}(1,2,...,p-1,0)$. For instance, if $p=2$ this representation $\mathfrak{g}\rightarrow\mathfrak{gl}_{2}$ is given by
$X\mapsto\begin{pmatrix}1&0\\ 0&1\end{pmatrix},~Y\mapsto\begin{pmatrix}0&a\\ 1&0\end{pmatrix},~Z\mapsto\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$.
Our restricted Lie algebra $\mathfrak{g}$ is solvable; it has a unique composition series of restricted subalgebras
$\mathfrak{g}\supset\langle X,Y\rangle\supset X\supset 0.$
Moreover, $\mathfrak{g}$ has no non-zero $p$-nilpotent elements.
Let $G:=G(\mathfrak{g})$ be the (unique) height 1 $k$-group associated to the Lie algebra $\mathfrak{g}$ (under the correspondence in [DG, II, §7, Prop. 4.1]). Let $N$ be the normal subgroup of $G$ associated to the ideal $\langle X,Y\rangle$ of $\mathfrak{g}$. Then $G$ is a non-nilpotent solvable infinitesimal $k$-group of order $p^{3}$, it has a unique composition series
$G\supset N\supset Z(G)\supset 1$
with successive quotients $\mu_{p}$, $\alpha_{p}$, and $\mu_{p}$ (taken in this order), and it contains no non-trivial unipotent subgroups. Observe that $G$ is $p$-reductive; and yet neither its normal subgroup $N$, nor its quotient $G/Z(G)$, are $p$-reductive.
4. Pseudo-$p$-reductivity
In this section we prove Theorem 1.8 and Corollary 1.9; in which we relate the notion of pseudo-$p$-reductivity with its smooth connected analogue.
Let $k$ be a field of characteristic $p>0$, and let $G$ be a smooth connected affine $k$-group.
Proof of Theorem 1.8.
Let $k^{\prime}/k$ be the minimal field of definition for $\mathscr{R}_{u}(G_{\overline{k}})$, $\pi^{\prime}:G_{k^{\prime}}\rightarrow G_{k^{\prime}}/\mathscr{R}_{u}(G_{k^{\prime}})=:G^{\prime}$ be the natural projection, and $i_{G}:G\rightarrow R_{k^{\prime}/k}(G^{\prime})$ the map associated to $\pi^{\prime}$ under the adjunction of extension of scalars and Weil restriction.
Proof.
We may assume without loss of generality that $k=k^{\operatorname{s}}$, since the formation of both $\mathscr{R}_{u}(G)$ and $\operatorname{Rad}_{u}(G_{1})$ commutes with base change by separable algebraic field extensions (by a standard argument of Galois descent).
We first prove the “only if” direction of (i) and (ii).
Suppose that $G_{1}$ is pseudo-$p$-reductive. Since $G_{1}\cap\ker i_{G}$ is a unipotent normal subgroup of $G_{1}$, it must be trivial. Hence $\operatorname{Lie}(\ker i_{G})$ is zero, and so $\ker i_{G}$ is étale. Now assume (for a contradiction) that $p=2$ and $G^{\prime}$ has a direct factor $H^{\prime}$ that is isomorphic to $\mathrm{SO}_{2n+1}$ for some $n\geq 1$. Consider the unipotent isogeny
$H^{\prime}\cong\mathrm{SO}_{2n+1}\rightarrow\operatorname{Sp}_{2n}$
described in [PY, §2] or [Va, §2]; its kernel $N^{\prime}$ is isomorphic to $(\alpha_{2})^{2n}$. Since $N^{\prime}$ is a normal subgroup of a direct factor of $G^{\prime}$, it is normal in $G^{\prime}$ itself.
Observe that $R_{k^{\prime}/k}(N^{\prime})$ is a unipotent normal subgroup of $R_{k^{\prime}/k}(G^{\prime})$, since the Weil restriction functor is continuous and it preserves unipotency. Consider the normal subgroup $\smash{V:=i_{G}^{-1}(R_{k^{\prime}/k}(N^{\prime}))}$ of $G$; it is unipotent as the property of unipotency is preserved by taking subgroups and group extensions. The Frobenius kernel $V_{1}$ of $V$ is a characteristic subgroup of $V$ by [Mi, 2.28]. Consequently, $V_{1}$ is a unipotent normal subgroup of $G_{1}$. Hence $V_{1}$ is trivial by assumption, and so $V$ is étale.
Since $i_{G}(G)$ is pseudo-split and pseudo-reductive, it contains a Levi subgroup $L$ by [CGP, Thm. 3.4.6]. By [CGP, Lem. 9.2.1] the quotient map $G\rightarrow i_{G}(G)$ induces an isomorphism on the respective maximal reductive $k^{\prime}$-quotients $\smash{G^{\prime}\xrightarrow{\sim}i_{G}(G)^{\prime}}$. So the inclusion $L\hookrightarrow i_{G}(G)$ induces an isomorphism $\smash{L_{k^{\prime}}\xrightarrow{\sim}G^{\prime}}$. Consequently, since $L$ is split and reductive, it has a direct factor $H$ that is isomorphic to $\mathrm{SO}_{2n+1}$. Let $N\cong(\alpha_{2})^{2n}$ be the largest unipotent normal subgroup of $H$. It follows that $R_{k^{\prime}/k}(N^{\prime})\cap L$ contains $N$ (one way to justify this is to appeal to [BMRS, Lem. 2.2]). But this contradicts the fact that $V$ is étale.
We next prove the “if” direction of (i) and (ii).
Suppose that $G_{1}$ is not pseudo-$p$-reductive. Recall from [Mi, 2.27(c)] that the formation of the Frobenius kernel commutes with base change by arbitrary field extensions. Since $\pi^{\prime}$ is smooth and surjective, it restricts to a (faithfully) flat homomorphism $\pi^{\prime}_{1}:(G_{1})_{k^{\prime}}\rightarrow G^{\prime}_{1}$ by [SGA3, VIIA, Prop. 8.2]; where $G^{\prime}_{1}$ denotes the Frobenius kernel of $G^{\prime}$.
Let $U$ be a non-trivial unipotent normal subgroup of $G_{1}$. Since $\pi^{\prime}_{1}$ is (faithfully) flat, $\pi^{\prime}(U_{k^{\prime}})=:U^{\prime}$ is a unipotent normal subgroup of $G^{\prime}_{1}$. There are two possibilities: either $U_{k^{\prime}}\subseteq\ker\pi^{\prime}$ or $U^{\prime}$ is non-trivial.
First consider the case where $U_{k^{\prime}}\subseteq\ker\pi^{\prime}=\mathscr{R}_{u}(G_{k^{\prime}})$. By construction $\ker i_{G}$ is the largest subgroup of $G$ satisfying the property that $(\ker i_{G})_{k^{\prime}}\subseteq\mathscr{R}_{u}(G_{k^{\prime}})$. So $U$ is contained in the unipotent normal subgroup $G_{1}\cap\ker i_{G}$ of $G_{1}$. Hence $\ker i_{G}$ is not étale.
Next consider the case where $U^{\prime}$ is non-trivial. Denote $\mathfrak{g}^{\prime}:=\operatorname{Lie}(G^{\prime})=\operatorname{Lie}(G^{\prime}_{1})$, and consider the ideal $\mathfrak{u}^{\prime}:=\operatorname{Lie}(U^{\prime})$ of $\mathfrak{g}^{\prime}$. Since $U^{\prime}$ is unipotent it intersects trivially with any maximal torus $T^{\prime}$ of $G^{\prime}$, hence $\mathfrak{u}^{\prime}\cap\operatorname{Lie}(T^{\prime})=0$. Then applying [Va, Lem. 2.1] tells us that $p=2$ and $G^{\prime}$ has a direct factor which is isomorphic to $\mathrm{SO}_{2n+1}$ for some $n\geq 1$. [N.B. We cannot simply deduce this from [Va, Thm. 2.2(a)] or [PY, Lem. 2.2] as the restricted unipotent radical is not necessarily a characteristic subgroup.]
It remains to prove the second assertion of (i). Suppose $p>2$.
Assume that $G$ is pseudo-reductive of minimal type. Since $p>2$, the root system of $G$ is reduced by [CGP, Thm. 2.3.10]. Hence $\ker i_{G}$ is central in $G$ by [CP1, Prop. 6.2.15]. By definition of minimal type, this implies $\ker i_{G}$ is trivial.
For the converse, assume that $G$ is perfect and $\ker i_{G}$ is étale. Since $G$ is pseudo-split and pseudo-semisimple, it decomposes as a commuting product $G=S_{1}...S_{r}$; where each $S_{j}$ is a pseudo-split (absolutely) pseudo-simple $k$-group. It follows that $\ker i_{G}=\ker i_{S_{1}}...\ker i_{S_{r}}$; in particular, $\ker i_{S_{j}}$ is étale for each j=1,,rj=1,...,r. We need the following result taken from [CP1, Ex. 7.1.1]: if $p>2$, $S$ is an absolutely pseudo-simple $k$-group and $\ker i_{S}$ is étale, then $\ker i_{S}$ is trivial. Applying this to our decomposition says that each $\ker i_{S_{j}}$ is trivial, hence so is $\ker i_{G}$. It follows that $G$ is pseudo-reductive of minimal type. This completes the proof. ∎
Proof of Corollary 1.9.
Proof.
Assume $G_{1}$ is pseudo-$p$-reductive. Applying Theorem 1.8 tells us that $\ker i_{G}$ is étale. Then $G$ is pseudo-reductive, as $\mathscr{R}_{u}(G)\subseteq\ker i_{G}$. Moreover, $\ker i_{G}$ is central in $G$ by [Mi, Rem. 12.39(a)]. The central quotient map $G\rightarrow i_{G}(G)$ induces a bijection on the respective root systems by [CGP, Prop. 2.3.15]. But $i_{G}(G)$ is pseudo-reductive with reduced $k^{\operatorname{s}}$-root system, so the root system of $G_{k^{\operatorname{s}}}$ is also reduced. ∎
Acknowledgements: The author would like to thank Michael Bate, Brian Conrad, Sean Cotner, Ben Martin, Timm Peerenboom, Gopal Prasad, Gerhard Röhrle and David Stewart, as well as an anonymous referee, for some helpful comments and discussions. Most of this research was funded by an Alexander von Humboldt postdoctoral grant, and currently the author is employed/funded by the University of Freiburg Mathematical Institute.